A car dealership has set a goal for its 15 salespeople of selling 150 new cars in an eight-week period. In the first six weeks of this period, an average of 19.5 cars were sold each week. How many cars must be sold in the remaining two weeks to achieve the 150 target?
Answer #1

Number of cars sold in first six weeks =Average number of cars sold each week*Number of weeks =19.5*6 =117

Number of cars to be sold in the remaining two weeks =Required target number of cars to be sold - Number of cars already sold =150 - 117 =33

Thus, 33 cars must be sold in the remaining two weeks to achieve the 150 target.